# Demystifying Misconceptions in Social Bots Research

### v3

@@
Research on social bots aims at advancing knowledge and providing solutions to one of the most debated forms of online manipulation. Yet, social bot research is plagued by widespread biases, hyped results, and misconceptions that set the stage for ambiguities, unrealistic expectations, and seemingly irreconcilable findings. Overcoming such issues is instrumental towards ensuring reliable solutions and reaffirming the validity of the scientific method. In this contribution, we review some recent results in social bots research, highlighting and revising factual errors as well as methodological and conceptual biases. More importantly, we demystify common misconceptions, addressing fundamental points on how social bots research is discussed. Our analysis surfaces the need to discuss research about online disinformation and manipulation in a rigorous, unbiased, and responsible way. This article bolsters such effort by identifying and refuting common fallacious arguments used by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research in the field.
@@ 1 Introduction
Human decision-making processes depend on the availability of high-quality information. In addition, a healthy society requires a shared understanding of issues and values to foster trust and cooperation. Misinformation and disinformation erode trust and emphasize divisions. Moreover, the presence of divergent or incompatible beliefs can make it difficult to reach a consensus and take effective collective action. The repercussions could be so severe as to delay responses to a deadly pandemic (Pierri et al., 2022) or endanger critical democratic processes (Pennycook et al., 2021).
@@
The science of misinformation seeks solutions to these problems (Lazer et al., 2018). However, this paramount endeavor can itself incur the same problems that it aims to overcome (West and Bergstrom, 2021; Altay et al., 2023). For example, scientific articles and publishers engage in a fierce competition for attention, similar to what is being experienced by news outlets. As a consequence, sensationalist claims and hyped results can sometimes be used as shortcuts to publication and other scientific accomplishments (West and Bergstrom, 2021). Then, confirmation bias is making its way into science in the form of citation bias: the preference for citing articles that support one’s results over those that challenge them (West and Bergstrom, 2021). And again, over-generalizations of scientific results and poor understanding of methodological and conceptual limitations give rise to multiple misconceptions about misinformation (Altay et al., 2023). These and other problems currently undermine the efficacy and credibility of our research efforts. Therefore, for the science of misinformation to benefit our society, we must first solve the problems within the science itself.
The present study concerns one of the many forms of online manipulation: social bots. In this context, let us make a brief clarification of what, for the purposes of this article, is a bot, a social bot, and a malicious social bot. In literature, the term bot is associated with several others that are often, and sometimes erroneously, used as synonyms. As discussed by Gorwa and Guilbeault (2020b), scholars, journalists, and policymakers often speak indiscriminately of bots, chatbots, botnets, sybils, and cyborgs, to cite a few. If we go back in time, the bots better known as chatbots referred to early programs that can converse with humans. Later in time, the word bot also came to mean a computer machine compromised, co-opted, and remotely controlled by malware (Yang et al., 2014). When these types of accounts work together, they form a network called botnet, whose methods of operation and proposals for detection have been studied by information security researchers (Yang et al., 2014; Antonakakis et al., 2017). With the advent of social media and social networks, bots were also framed as compromised network accounts, called sybils in the information security field. Considering the bots that populate online social networks—conventionally called social bots (Ferrara et al., 2016)—there are works that categorize such accounts according their intent or their capacity to imitate humans (Stieglitz et al., 2018), as in the case of social bots promoting or suppressing certain political narratives (Pratelli et al., 2024) or facilitating the spread of disinformation (Shao et al., 2018). Others differentiated social bots according to their activity on the platform, identifying those that systematically retweet the posts of political candidates (Cresci et al., 2017), or those that mass-follow a brand or character to boost their popularity (Cresci et al., 2015). There even exist social bots with no malicious intent that rather serve a useful purpose for those who follow them (Cresci, 2020; Savage et al., 2016). For example, these bots can act as news aggregators (Lokot and Diakopoulos, 2016).
Among the many diverse existing definitions, here we define bots as social media accounts that are totally or partially automated. By focusing on the use of automation, we adopt a definition that is both widely used and based on an objective criterion (Ferrara et al., 2016; Wischnewski et al., 2021; Ferrara, 2022). This is in contrast to those based on the intent (malicious, neutral, or benign), on the capacity to imitate humans, or on the type of activity carried out, which are inherently more subjective. Nonetheless, we clarify that independently of the chosen definition, the issues highlighted in the remainder of this article as well as the proposed solutions, are broadly valid and applicable, ensuring that adhering to a different definition does not diminish the overall validity of the analyses and discussions herein.
@@
Thanks to their automated nature, ease of use, and negligible marginal cost, social bots can be employed as expendable tools for spreading problematic content. Understanding the role and activity of bots in large-scale manipulation campaigns is important, as it can inform strategies to curb misinformation and disinformation (Shao et al., 2018). For this reason, social bots have attracted considerable scholarly and media attention (Allem and Ferrara, 2018; Assenmacher et al., 2020; González-Bailón and De Domenico, 2021; Chen et al., 2021). One of the latest such episodes was the public dispute on the real number of bots on Twitter that took place before Elon Musk’s acquisition of the platform (Varol, 2023). However, despite many years of research, the science of social bots is replete with the same problems that also plague the science of misinformation. These originate from uncorrected factual errors, from the presence of biases in how scientific results are cited and discussed, and from a wide array of methodological and conceptual issues that set the stage for ambiguities, misunderstandings, unrealistic expectations, as well as contrasting and seemingly irreconcilable findings (Rauchfleisch and Kaiser, 2020; Hays et al., 2023).
In our present analyses, we focus on some recently published articles about social bots. A thorough analysis of their methods, assumptions, results, and conclusions reveals several pitfalls ranging from factual inaccuracies and methodological flaws to more subtle incidents such as misrepresentation of the existing literature and conceptual errors. Our discussion links these issues to common misconceptions in social bots and misinformation research. Our critical analysis thus serves a twofold goal. On the one hand, we highlight and revise factual errors, methodological biases, and conceptual issues. On the other hand, demystifying common misconceptions allows us to address fundamental issues on how science is produced and discussed (West and Bergstrom, 2021). Overcoming such issues is instrumental towards ensuring the credibility of science.
@@ 2 Methodological issues
Social bot detection is a critical and increasingly challenging task in the realm of online safety and cybersecurity (Cresci, 2020). In practice, it often involves the use of machine learning algorithms in a binary classification setting, with the goal of distinguishing human-operated accounts from automated ones—the social bots—on various online platforms. The machine learning models are trained on features extracted from user profiles, behaviors, network interactions, and linguistic patterns acquired from user posts with the goal of detecting subtle differences between authentic users and automated entities. Methodological issues in social bots research include a mix of typical machine learning pitfalls and unique issues inherent in the evolving nature of online social ecosystems, which we exemplify in the following sections.
@@
Model drift happens when a model relies on a set of features that sensibly diverge between train and test samples. Model drift generally happens when data slowly change distribution over time. However, it can also happen when the training and prediction datasets are significantly different. Generally, model drift can slowly reduce model accuracy over time, but it can also misleadingly show that the trained model performs well due to the exploitation of feature peculiarities not found in the prediction dataset. In bot detection, the problem may surface when knowledge of some peculiar characteristic of the accounts is known beforehand and exploited at training time, allowing for near-perfect detection performance. For example, each botnet typically exhibits some peculiar characteristics that surface as a result of how the bots were created, their goals, or any other shared characteristic between them (Zhang et al., 2016; Mazza et al., 2022). These peculiar characteristics might set the bots apart from the average behavior of human-operated accounts (Cresci et al., 2020). However, most of the time, such peculiar characteristics are known only for botnets that have already been identified and are instead unknown for still undetected botnets. Here it is important to recall that the end goal of any bot detector—as with any machine learning classifier—is to accurately detect unknown bots from a set of never-seen-before data. Thus, the detection performance on a held-out portion of known data serves merely as a proxy for the real-world detection performance in-the-wild. However, if a bot detector exploits known characteristics at training time for classifying instances of known bots, it might learn correlations that do not generalize to the unknown bots that the classifier will be tasked to detect later on.
Another way in which a model drift occurs in bot detection is when a detector exploits knowledge of how the accounts were collected—be them the malicious bots or the genuine ones—as a means to predict their class (i.e, automated or otherwise). The correlation between the information on how certain accounts were collected and their class is spurious with respect to the task of bot detection, which makes features based on such information dangerously misleading. Consequently, a detector exploiting these features would achieve excellent performance on known data but would likely exhibit poor capacity to generalize to new data with different characteristics.
@@
Such situation can lead to the proposal of ostensibly superior models, which however represent a regression rather than advancement in addressing the bot detection problem. In addition, the same models could also be exploited to unduly undermine the performance of other models. This was recently done by Gallwitz and Kreil (2022) and Hays et al. (2023), who trained simple but unrealistic bot detectors using features tightly bound to the training dataset, such as whether an account was verified by Twitter—which was among the features used to select and label the data—to exaggerate the limitations of more general state-of-the-art models. The use of a feature used to select and label the data can also be framed as a feature leakage issue: it is crucial to refrain from using a feature that is utilized to select and label the data as a training input for the model since it will eventually lead to unreliable and inaccurate predictions.
@@
Another challenge in bot detection arises from the multitude of disparate definitions, detectors, and benchmark datasets (Cresci, 2020). For example, a critical issue in evaluating bot detectors is the possible inconsistency between the bot definition assumed by the detector and that used by the evaluator. This inconsistency can lead to biased and unfair evaluations as it is conceptually flawed to evaluate a detector using a different definition than that employed by the detector itself. A tool trained to identify a specific type of bot should be assessed based on its ability to detect that specific type, not alternative ones for which it was not designed. This misalignment in definitions can result in evaluations that misrepresent the detector’s true performance and capabilities. Therefore, while it is perfectly acceptable to discuss—and even criticize—the social bot definition used by a detector, evaluators must ensure that the definition used to evaluate a detector aligns with the detector’s, or be very clear about the differences and their implications towards the results of the evaluation. We point out, however, that there exists a trade-off between the need for consistency in using the same definition as the detector and the necessity of assessing the detector’s generalizability, which requires testing it against bots with different characteristics (Cresci, 2020). Striking this balance depends on multiple factors, including the practical context in which the detector is used and whether the detector is specialized towards specific types of bots or general-purpose.
@@
The evaluation landscape is further complicated by the seemingly excellent performance achieved by models that overfit to specific evaluation datasets. [^1] We refer to overfitting when a model depends too much on irrelevant features of the training data, with the result that it performs well on the training data but generalizes poorly to unseen data (Bramer, 2007). The issue arises from the inappropriate comparison of a general-purpose model with an extremely specialized—potentially overfitted—one. As an example, Hays et al. (2023) selected some datasets to train trivial classifiers with as few as four features. Then, they compared the trivial classifier to more complex, general, state-of-the-art models, on the same dataset where the trivial model had been trained, suggesting that very few features are sufficient to identify bots in a given dataset. While this approach can highlight biases in certain datasets, the comparison itself is unfair and should not be used to undermine the utility of general-purpose detectors or to criticize their performance.
In addition to carefully choosing the models with which to compare performance, attention must also be devoted to the data used for the evaluation. As an example of problematic use of evaluation data, Gallwitz and Kreil (2022) selected prominent public accounts, such as members of the US Congress and Nobel Prize Laureates, to challenge the accuracy of a widely-used bot detection tool. These institutional accounts are usually managed by a team of social media managers via dedicated software, and not directly operated by the corresponding public figures. In other words, this dataset is a non-representative sample of the accounts that inhabit social platforms, and evaluating a general-purpose bot detector solely on these accounts likely leads to an under- or overestimation of the detector’s performance. Although relevant towards evaluating the performance of the detector in a very specific context, this evaluation does not warrant a generalized approval or criticism of the tool. More broadly, since bot detection methods—whether based on machine learning or on human annotation—have inherent error margins, accuracy results can be manipulated by carefully choosing test examples.
@@
The issue of cherry-picking can manifest in unfair comparisons, but extends well beyond it. A concerning trend involves selectively including, excluding, or misrepresenting prior literature to propose a narrative that aligns favorably with one’s own findings. Recently, this has been done to cast seemingly new criticism against some bot detectors (Hays et al., 2023). These claims can only be made by omitting a significant amount of literature (Varol et al., 2017; Yang et al., 2019, 2020; Sayyadiharikandeh et al., 2020; Cresci, 2020). On the other hand, Gallwitz and Kreil (2022) cite their own unpublished results—missing the vetting of peer review and shared via blog posts and presentations—multiple times, to give a false impression of prior research supporting their claims. Such selective referencing can not only distort the perceived reliability and novelty of a work, but also skew the representation of the state-of-the-art in bot detection. While citing unpublished work is not always bad per se, by self-citing under different forms multiple unvetted claims about the results of a single analysis, authors may create an illusion of authority. And by omitting works that have already made certain contributions or conclusions, authors may create an illusion of originality, potentially overshadowing a substantial body of pre-existing research that has contributed to the same or comparable insights. This not only undermines the integrity of the academic discourse but also detracts from the collective acknowledgment and recognition owed to the broader community of researchers who have previously advanced the field.
Addressing the cherry-picking issue in social bots research necessitates a reliance on expert reviewers deeply versed in the nuances of the field, capable of discerning the strategic omission of relevant literature, definitions, methods, or data. Paradoxically, the escalating trend of publications (Haghani et al., 2022), particularly in hot topics like bot detection (Cresci, 2020), poses a formidable challenge, as the growing demand for reviewers outpaces the available pool of experts. This discrepancy surfaces a crucial tension in maintaining the quality and rigor of peer review processes within rapidly evolving research landscapes (Van Noorden, 2023). The field is now confronted not only with the difficulty of establishing a definitive benchmark, but also with the challenge of navigating through a heterogeneous literature landscape, where discerning genuine advancements becomes a complex task amid the noise produced by contributions of varying quality.
@@
The straw-man fallacy consists of misrepresenting someone else’s research and then criticizing the misrepresentation instead of the original research. For example, Gallwitz and Kreil (2022) make a generic and sensationalist criticism against all social bots research based on a misrepresentation of the literature. First, they claimed that «in our impression, most of the study authors we contacted were fully aware that the lion’s share or even all of the accounts they counted as ‘bots’ were, in fact, humans.» Such strong insinuation that the majority of researchers in the field knowingly misrepresented human accounts as bots is a serious allegation that requires robust evidence. However, this accusation is only based on the authors’ impression, lacking any scientific validity. Second, they claimed that a systematic evaluation of certain bot predictions in real-world scenarios have never been done prior to their work (Gallwitz and Kreil, 2022), pointing to a lack of manual validation and publicly available datasets for such evaluations. However, this claim is largely exaggerated. The majority existing bot detectors have been rigorously evaluated on training and test datasets, with many of these datasets being publicly available. While it is indeed important to manually validate the labels assigned by detectors and share this validation data, the generalized criticism that no systematic evaluation had been done before is unwarranted and potentially misleading, as such statement overlooks the extensive evaluations that have already been conducted in the field. Another problematic claim is that about Gallwitz and Kreil’s failure to reproduce an experiment by Keller and Klinger (2019). Here, they applied a different bot detection model than the one used by Keller and Klinger, and then attributed the changes in bot score distributions to inconsistencies in Keller and Klinger’s detector, without mentioning that such differences could instead be due to the different choices they made with respect to Keller and Klinger.
@@
A further manifestation of the straw-man methodology is the claim that bot detection is exclusively a supervised machine learning task. This wrong assumption could perhaps be explained by the fact that supervised machine learning is the traditional way in which the task was tackled, and by the multitude of existing supervised bot detectors (Cresci, 2020). Despite their prevalence, however, a robust body of literature highlighted that supervised detectors suffer from lack of generalizability and transferability, among other issues (Cresci, 2020; Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021). A variety of unsupervised and semi-supervised bot detectors were proposed as possible solutions to these problems, considering groups instead of single accounts and studying features like the temporal patterns of their actions or the structures of their graph representations (Cresci, 2020). Evaluating the generalization capabilities of bot detectors is precisely the goal of Hays et al. (2023). However, they chose to do so by adopting a supervised learning methodology that has been proven to suffer from critical issues when it comes to generalizability and transferability. Another issue with their methodology is that it does not fully support their conclusions. Specifically, Hays et al. leverage their results to criticize the generalization capabilities of all bot detectors, despite having experimented only with a narrow set of supervised methods. Similarly, Gallwitz and Kreil conclude that "the field of social bot research is fundamentally flawed", despite having investigated only a tiny fraction of the research based on one single bot detector. Neither of the two aforementioned works make any distinction between the different types of bot detectors proposed to date—including differences between supervised ones based on the analysis of individual accounts vs. unsupervised ones based on the analysis of groups of accounts (Cresci, 2020). In (Hays et al., 2023), this is a textbook example of circular reasoning: the goal of the analysis is to evaluate whether current bot detectors are generalizable. However, the analysis only evaluates supervised detectors, which have known generalization issues (Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021). The unsurprising—and biased—result is that the considered detectors fail to generalize, which is used to criticize the whole field. Future work should avoid oversimplifying social bot detection as a solely supervised task. Instead, all works that propose new detectors, that use already existing ones, or that discuss the state-of-the-art, should acknowledge the multitude of approaches to the task, together with their strengths and weaknesses. Then, the type of bot detector to develop, use, or discuss in a given study should be chosen so as to be adequate and consistent with the objective of the study.
@@
In machine learning and data science tasks, high-quality data is the linchpin for robust model development and insightful analyses (Halevy et al., 2009). As mentioned above, numerous datasets of bot and human accounts have been published over the years. On the one hand, this made it easier to thoroughly test the performance of new bot detectors. On the other hand, it introduced the risk of cherry-picking favorable datasets, potentially distorting reported performances and introducing bias. Furthermore, the temporal heterogeneity stemming from the availability of datasets spanning varying time periods—ranging from recent to decade-old—can pose additional challenges when detectors are trained on outdated data, undermining the relevance of their performance in contemporary settings. When considering the representativeness of published benchmark datasets with respect to the bots that currently inhabit online platforms, the readers should be aware that the accounts included in a dataset are likely a few years older than the publication date of the dataset itself. Moreover, some datasets are built by combining or extending previously published resources. For example, the pan-2019 dataset (Rangel and Rosso, 2019) is largely based on other datasets published between 2011 and 2017. Hence, the publication date of a dataset should be considered as a generous upper bound of the recency of the accounts therein. Even under this relaxed assumption, in consideration of the rapid pace of change of the landscape of online harms and of the availability of more comprehensive and more recent datasets (Feng et al., 2021), we may run the risk of capturing datasets that are no longer representative of the actual state of the platforms. This is particularly troublesome in light of the known evolutionary behavior of social bots, typical of adversarial settings, which requires constant updates of data and methods (Cresci et al., 2021). However, the majority of newly published bot detectors are at least partially based on dated—if not outright obsolete—training data. Obsolete datasets can be used to criticize other people’s detection methods or to promote one’s own method as better than the state of the art.

### v4

@@
Research on social bots aims at advancing knowledge and providing solutions to one of the most debated forms of online manipulation. Yet, social bot research is plagued by widespread biases, hyped results, and misconceptions that set the stage for ambiguities, unrealistic expectations, and seemingly irreconcilable findings. Overcoming such issues is instrumental towards ensuring reliable solutions and reaffirming the validity of the scientific method. Here, we discuss a broad set of consequential methodological and conceptual issues that affect current social bots research, illustrating each with examples drawn from recent studies. More importantly, we demystify common misconceptions, addressing fundamental points on how social bots research is discussed. Our analysis surfaces the need to discuss research about online disinformation and manipulation in a rigorous, unbiased, and responsible way. This article bolsters such effort by identifying and refuting common fallacious arguments used by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research.
@@ 1 Introduction
Human decision-making processes depend on the availability of high-quality information and a healthy society requires a shared understanding of issues and values. Misinformation and disinformation erode trust and emphasize divisions. Moreover, the presence of divergent or incompatible beliefs can hinder reaching consensus and taking effective collective action. The repercussions could be so severe as to delay responses to a deadly pandemic (Pierri et al., 2022) or endanger democratic processes (Pennycook et al., 2021).
@@
The science of misinformation seeks solutions to these problems (Lazer et al., 2018). However, this paramount endeavor can itself incur the same problems that it aims to overcome (West and Bergstrom, 2021; Altay et al., 2023). For example, scientific articles and publishers engage in a fierce competition for attention, much like mainstream news outlets. As a consequence, sensationalist claims and hyped results are sometimes used as shortcuts to publication and scientific recognition (West and Bergstrom, 2021). Confirmation bias is also making its way into science, in the form of citation bias: the preference for citing articles that support one’s results over those that challenge them (West and Bergstrom, 2021). And again, over-generalizations of scientific results and poor understanding of methodological and conceptual limitations give rise to multiple misconceptions about misinformation (Altay et al., 2023). These and other problems currently undermine the efficacy and credibility of our research efforts. Therefore, for the science of misinformation to benefit our society, we must first solve the problems within the science itself.
The present study concerns one of the many forms of online manipulation: social bots. Among the many diverse existing definitions (Gorwa and Guilbeault, 2020), here we define social bots as social media accounts that are totally or partially automated. Hereafter, we use the terms “social bot” and “bot” interchangeably, always in line with the aforementioned definition.
@@
Due to their automation, simplicity, and low operating cost, social bots can be easily used as expendable tools for spreading problematic content. Understanding the role and activity of bots in large-scale manipulation campaigns is important, as it can inform strategies to curb mis- and disinformation (Shao et al., 2018). For this reason, social bots have attracted considerable scholarly and media attention (Allem and Ferrara, 2018; Assenmacher et al., 2020; González-Bailón and De Domenico, 2021; Chen et al., 2021). A recent example is the public dispute over Twitter’s bot count preceding Elon Musk’s acquisition (Varol, 2023). However, despite many years of research, the science of social bots is replete with the same problems that also plague the science of misinformation. These originate from uncorrected biases in how scientific results are cited and discussed, and a wide array of methodological and conceptual issues that set the stage for ambiguities, misunderstandings, unrealistic expectations, as well as contrasting and seemingly irreconcilable findings (Rauchfleisch and Kaiser, 2020; Hays et al., 2023).
Here, we develop a critical and theoretically grounded perspective on key methodological and conceptual challenges in social bots research. To this end, we discuss a large set of consequential issues for the field, based on their recurrence, impact, and potential for misunderstanding. In discussing each issue, we refer to a small number of recent studies that serve as illustrative examples of the broader problems. Therefore, while this article does not aim to be exhaustive in terms of the analyzed literature, the range of issues addressed here is extensive and comprehensive—as reflected in Table LABEL:tab:checklist. Our critical analysis thus serves a twofold goal. On the one hand, we highlight and revise methodological biases and conceptual issues in social bots research. On the other hand, demystifying common misconceptions allows us to address fundamental issues on how science is produced and discussed (West and Bergstrom, 2021). Overcoming such issues is instrumental in ensuring the credibility of science.
@@ 2 Methodological issues
In this section, we focus on issues that arise from how social bots research is conducted and evaluated, covering aspects such as data selection, model training, and the dissemination of results.
Social bot detection is a challenging task in the realm of online safety and cybersecurity (Cresci, 2020). In practice, it often involves the use of machine learning algorithms in a binary classification setting, with the goal of distinguishing human-operated accounts from automated ones—the social bots. The machine learning models are trained on features extracted from user profiles, behaviors, network interactions, and linguistic patterns acquired from user posts, with the goal of detecting subtle differences between authentic users and automated entities. Methodological issues in social bots research include a mix of typical machine learning pitfalls and unique issues inherent in the evolving nature of online social ecosystems, which we exemplify in the following sections.
@@
Model drift arises when the features a model relies on differ between training and prediction data—whether due to gradual shifts over time or inherent dataset differences. It can erode accuracy or falsely suggest strong performance by exploiting training-specific patterns absent at prediction time. In bot detection, the problem may surface when knowledge of some peculiar characteristic of the accounts is known beforehand and exploited at training time, allowing for near-perfect detection performance. For example, each botnet typically exhibits some peculiar characteristics resulting from the bot creation process, their goals, or any other shared characteristic (Zhang et al., 2016; Mazza et al., 2022). These peculiar characteristics might set the bots apart from the average behavior of human-operated accounts (Cresci et al., 2020). However, typically such characteristics are known only for botnets that have already been identified and are instead unknown for still undetected botnets. Here we recall that a classifier’s detection performance on a held-out portion of known data serves merely as a proxy for its real-world detection performance in-the-wild. In fact, the end goal of any bot detector is to accurately detect unknown bots from a set of never-seen-before data. However, if a bot detector exploits known characteristics at training time for classifying instances of known bots, it might learn correlations that do not generalize to the unknown bots that the classifier will be tasked to detect later on.
@@
A specific instance of model drift in bot detection is when a detector exploits knowledge of how the accounts were collected—be them the malicious bots or the genuine ones—as a means to predict their class (i.e, automated or otherwise). The correlation between the information on how certain accounts were collected and their class is spurious with respect to the task of bot detection, which makes features based on such information dangerously misleading. Consequently, a detector exploiting these features would achieve excellent performance on known data but would exhibit poor capacity to generalize to new data with different characteristics. Leveraging features used to select and label the data can also be framed as a feature leakage issue, since those features would unfairly advantage the detector. Such a situation can lead to the proposal of ostensibly superior models, which, however, represent a regression rather than an advancement in addressing the bot detection problem. In addition, the same models could also be exploited to unduly undermine the performance of other models. For example, some scholars trained simple but unrealistic bot detectors using features tightly bound to the training dataset, such as whether an account was verified by Twitter—which was among the features used to select and label the data—to exaggerate the limitations of more general state-of-the-art models (Gallwitz and Kreil, 2022; Hays et al., 2023).
@@
Other challenges arise from the multitude of disparate definitions, detectors, and benchmark datasets (Cresci, 2020). For example, a critical issue in evaluating bot detectors is the possible inconsistency between the bot definition assumed by the detector and that used by the evaluator. This inconsistency can lead to biased and unfair evaluations, as it is conceptually flawed to evaluate a detector using a different definition than that employed by the detector itself. A tool trained to identify a specific type of bot should be assessed based on its ability to detect that specific type, not alternative ones for which it was not designed. This misalignment in definitions can result in evaluations that misrepresent the detector’s true performance and capabilities. Therefore, while it is perfectly acceptable to discuss—and even criticize—the social bot definition used by a detector, evaluators must ensure that the definition used to evaluate a detector aligns with the detector’s, or be very clear about the differences and their implications towards the results of the evaluation. However, there exists an important trade-off between the need for consistency in using definitions and the necessity of assessing the detector’s generalizability, which requires testing it against bots with different characteristics (Cresci, 2020). Striking this balance depends on multiple factors, including the practical context in which the detector is used and whether the detector is specialized towards specific types of bots or general-purpose.
@@
The evaluation landscape is further complicated by the seemingly excellent performance achieved by models that overfit to specific evaluation datasets. [^1] We refer to overfitting when a model depends too much on irrelevant features of the training data, with the result that it performs well on the training data but generalizes poorly to unseen data (Bramer, 2007). The issue arises from the inappropriate comparison of a general-purpose model with an extremely specialized—potentially overfitted—one. As an example, Hays et al. (Hays et al., 2023) selected some datasets to train trivial classifiers with a very small number of features. Then, they compared the trivial classifiers to more complex, general, state-of-the-art models, on the same dataset where the trivial models had been trained, suggesting that very few features are sufficient to identify bots. While this approach can highlight biases in certain datasets, the comparison itself is unfair and should not be used to undermine the utility of general-purpose detectors or to criticize their performance.
Attention should also be devoted to the data used for the evaluation. As an example of problematic use of evaluation data, Gallwitz & Kreil (Gallwitz and Kreil, 2022) challenged the accuracy of a widely-used bot detector based on its results on a small dataset of public figures’ accounts. However, these accounts—typically managed by social media teams—do not represent the broader platform user base, possibly leading to under- or overestimation of the detector’s performance. While useful in a specific context, such evaluation does not justify generalized conclusions about the detector’s effectiveness. More broadly, since bot detection methods—whether based on machine learning or on human annotation—have inherent error margins, accuracy results can be manipulated by carefully choosing test examples.
@@
The issue of cherry-picking extends beyond unfair comparisons. A concerning trend involves selectively including, excluding, or misrepresenting prior literature to propose a narrative that aligns favorably with one’s own findings. Recently, this has been done to cast seemingly new criticism against certain bot detectors (Hays et al., 2023). However, the impression of novelty in such criticism can only be made by omitting a significant amount of literature (Varol et al., 2017; Yang et al., 2019, 2020; Sayyadiharikandeh et al., 2020; Cresci, 2020). An opposite—but equally misleading—practice involves repeatedly citing one’s own unpublished results, which could give a false impression of prior research supporting one’s claims (Gallwitz and Kreil, 2022). Selective referencing can not only distort the perceived reliability and novelty of a work, but also skew the representation of the state-of-the-art. While citing unpublished work is not always bad per se, by self-citing under different forms multiple unvetted claims about the results of a single analysis, authors may create an illusion of authority. And by omitting works that have already made certain contributions or conclusions, authors may create an illusion of originality, potentially overshadowing a substantial body of pre-existing research that has contributed comparable insights. This not only undermines the integrity of the academic discourse but also detracts from the collective acknowledgment and recognition owed to the broader community of scholars who have previously advanced the field.
Addressing the cherry-picking issue necessitates a reliance on expert reviewers deeply versed in the nuances of the field, capable of discerning the strategic omission of relevant literature, definitions, methods, or data. Paradoxically, the escalating trend of publications (Haghani et al., 2022), particularly in hot topics like bot detection (Cresci, 2020), poses a formidable challenge, as the growing demand for reviewers outpaces the available pool of experts. This discrepancy surfaces a crucial tension in maintaining the quality and rigor of peer review processes within rapidly evolving research landscapes (Van Noorden, 2023). The field is now confronted not only with the difficulty of establishing a definitive benchmark, but also with the challenge of navigating through a heterogeneous literature landscape, where discerning genuine advancements becomes a complex task amid the noise produced by contributions of varying quality.
@@
The straw-man fallacy consists of misrepresenting someone else’s research and then criticizing the misrepresentation instead of the original research. A common manifestation is the claim that bot detection is exclusively a supervised machine learning task. This wrong assumption could perhaps be explained by the fact that supervised machine learning is the traditional way in which the task was tackled, and by the multitude of existing supervised bot detectors (Cresci, 2020). Despite their prevalence, however, a robust body of literature highlighted that supervised detectors suffer from lack of generalizability and transferability (Cresci, 2020; Echeverr´ıa et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021). A variety of unsupervised and semi-supervised detectors were proposed as possible solutions to these problems, considering groups instead of single accounts and studying features like the temporal patterns of their actions or the structures of their graph representations (Cresci, 2020). Nonetheless, some studies specifically focus on evaluating the generalization capabilities of bot detectors, but only consider supervised detectors with known generalization deficiencies. The unsurprising result is that the considered detectors fail to generalize, which is used to criticize the whole field—a textbook example of circular reasoning. For instance, Hays et al. (Hays et al., 2023) criticize the generalization capabilities of all bot detectors, despite having experimented only with a narrow set of supervised methods. Similarly, Gallwitz & Kreil (Gallwitz and Kreil, 2022) conclude that «the field of social bot research is fundamentally flawed», despite having investigated only a tiny fraction of the research based on one single bot detector. The issue lies in the broad conclusions of such studies, which are unsupported by their methodology and results. Neither of the two aforementioned works made distinctions between the different types of bot detectors proposed to date. A further example is the claim that a systematic evaluation of bot predictions in real-world scenarios have never been done (Gallwitz and Kreil, 2022). The accusation points to a lack of manual validation and publicly available datasets. However, the majority of existing bot detectors have indeed been evaluated on training and test datasets, with many of these datasets being publicly available. While we acknowledge that such evaluations may have inherent limitations, and that it is important to manually validate bot labels and to share this validation data, the generic criticism that no systematic evaluation has been done before is unwarranted and potentially misleading.
Therefore, future work should avoid oversimplifying social bot detection as a solely supervised task. Instead, works that propose new detectors, that use already existing ones, or that discuss the state-of-the-art, should acknowledge the multitude of approaches to the task, with their strengths and weaknesses. Then, the type of evaluation, as well as the type of bot detector to develop, use, or discuss in a given study should be chosen so as to be adequate and consistent with the objective of the study.
@@
In machine learning and data science, high-quality data is the linchpin for robust model development and insightful analyses (Halevy et al., 2009). As mentioned above, numerous datasets of bot and human accounts have been published over the years. On the one hand, this made it easier to thoroughly test the performance of new bot detectors. On the other hand, it introduced the risk of cherry-picking favorable datasets, potentially distorting reported performances and introducing bias. Furthermore, the temporal heterogeneity stemming from the availability of datasets spanning varying time periods—ranging from recent to decade-old—can pose additional challenges when detectors are trained on outdated data, undermining the relevance of their performance in contemporary settings. When considering the representativeness of published benchmark datasets with respect to the bots that currently inhabit online platforms, the readers should be aware that the accounts included in a dataset are likely a few years older than the publication date of the dataset itself. Hence, the publication date of a dataset should be considered as a generous upper bound of the recency of the accounts therein. Even under this relaxed assumption, in consideration of the rapidly evolving landscape of online harms and the availability of more comprehensive and recent datasets (Feng et al., 2021), we risk relying on datasets that are no longer representative of the actual state of the platforms. This is particularly troublesome in light of the known evolutionary behavior of social bots, typical of adversarial settings, which requires constant updates of data and methods (Cresci et al., 2021). However, many newly published bot detectors are at least partially based on obsolete training data.